OpenStudy (anonymous):

Need some help please.. Obtain the general solution: y''+y=12(cos x)^2

OpenStudy (anonymous):

Problem shows as: \[y''+y=12\cos^2x\] I found my Yc, but i need help finding Yp please

OpenStudy (anonymous):

\[Yc=C1cosx+C2sinx\]

OpenStudy (jtvatsim):

I'm thinking maybe a way to start is to integrate cos^2(x) twice... I'm still trying to work it out.

OpenStudy (anonymous):

I was wondering if maybe i need to use a trig identity..

OpenStudy (jtvatsim):

maybe we can start by recognizing that \[\cos^2 x = \frac{1}{2}(1 + \cos(2x))\] so the question is essentially dealing with \[\cos(2x)\]

OpenStudy (jtvatsim):

phew, I think I found it.

OpenStudy (anonymous):

ahhhhhh.... i see, ok so would Yp=A+B(cos2x)?

OpenStudy (jtvatsim):

that should probably work, I used more terms because I wasn't sure what I was going to need, but if you get something like Yp = -2 cos 2x + 6 as your final answer, then your choice worked.

OpenStudy (jtvatsim):

yeah, your choice works also (and is much faster btw). :)

OpenStudy (anonymous):

Awesome! Thanks for your help! @jtvatsim

OpenStudy (jtvatsim):

not a problem, you hit the nail on the head when you were thinking of trig identities... you deserve a medal yourself. :)

OpenStudy (anonymous):

Right on! My first one lol @jtvatsim

OpenStudy (jtvatsim):

It's always nice to get medals on your own questions... doesn't happen too often for me. :)